OpenStudy (anonymous):

In 2000 the population of Austraiia was 19.0 M, in 2010 it was 22.6 M. The exponential model P=P_0 e^kt describes the Australian population, P, in millions, t years after 2000.

OpenStudy (anonymous):

Find the growth rate, k, to 5 decimal places.

OpenStudy (phi):

t is the number of years since 2000. i.e. for year 2000, t=0 put in t=0 and P= 19.0 to create one equation then do the same for t= 10 (meaning 2010) and P= 22.6 what two equations do you get ?

OpenStudy (anonymous):

\[19.0=19.0_0 e ^{k+0}\]

OpenStudy (phi):

almost. the \( 19.0_0 \) should stay \(P_0\) also, it is k*t not k+t

OpenStudy (anonymous):

oh yea my bad with the k+0 thing why does it stay as \[p _{0}\]

OpenStudy (phi):

P0 and k are the unknowns. However, we are given P (population) and time t that means we can form 2 equations (with 2 unknowns) can you write the 2 equations?

OpenStudy (anonymous):

Well I'm assuming that \[19.0=P _{0} e ^{k*0}\] is one but im not sure of the other @phi

OpenStudy (phi):

the other one has a different P and a different t in 2010 it was 22.6 M t is the years since 2000

OpenStudy (anonymous):

oh alright the other one is

OpenStudy (anonymous):

\[22.6=P _{0} e ^{k*10}\]

OpenStudy (phi):

ok. notice in the first one you have an exponent k*0 which is just 0 also e^0 is 1 (anything to the 0 power is 1) so the first equation simplifies to 19.0= P0 that means we have found P0 use that value in the 2nd equation

OpenStudy (anonymous):

22.6=19.0^k(10)

OpenStudy (phi):

ok, except you left out the e \[ 22.6= 19.0 e^{10k} \]

OpenStudy (phi):

now divide both sides by 19.0 \[ \frac{22.6}{19.0} = e^{10k} \] now take the "natural log" of both sides. can you do that ?

OpenStudy (anonymous):

\[\frac{ \ln(22.6) }{ \ln(19.0) }=10?\]

OpenStudy (anonymous):

10k*

OpenStudy (phi):

you take the ln of both sides like this \[ \ln\left( \frac{22.6}{19.0}\right) = \ln \left( e^{kt}\right) \]

OpenStudy (anonymous):

does the 10 and the k not come down then?

OpenStudy (phi):

which is also \[ \ln\left( \frac{22.6}{19.0}\right) = 10k \] yes the 10k "comes down" (It's important to memorize this property)

OpenStudy (anonymous):

so then I just get the ln of that and solve

OpenStudy (phi):

* kt should have been 10k yes, we get (after dividing both sides by 10) \[ 0.1\cdot \ln\left( \frac{22.6}{19.0}\right) = k \]

OpenStudy (anonymous):

0.0173511?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes. you can try it in the equation \[ P = 19 e^{0.0173511 t} \]

OpenStudy (phi):

the only tweak is they want k to five decimal places 0.01735